OpenStudy (anonymous):

lim as x->ln2 of ((e^3x)-8)/((e^2x)-4)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \ln2}(e ^{3x}-8)/(e ^{2x}-4)\] \[e ^{lnx}=x\] and \[3\ln x = (\ln x)^3\] that should help you get to the answer. Let me know if you need more help

OpenStudy (anonymous):

And make use of De L'Hôpital.

OpenStudy (badhi):

I would use the substitution, $$y=e^x$$ then, $$x\to \ln 2 \implies y\to 2$$ $$\begin{align*} \lim\limits_{x\to \ln2} \frac{e^{3x}-8}{e^{2x}-4}&=\lim\limits_{y\to2}\frac{y^3-8}{y^2-4} \end{align*}$$

OpenStudy (anonymous):

Thanks everyone! Azolator, I almost got it with your hints but BAdhi's method of substitution was so slick.

OpenStudy (anonymous):

I just want to point out that BAdhi's substitution worked because of those properties so make sure you know why you can substitute some things and not others